OpenStudy (anonymous):

y=-x^2-6x I need to find vertex and axis of symmetry. I got 3,9 for vertex?

OpenStudy (anonymous):

no. vertex is (-3, y)

OpenStudy (anonymous):

Is this how I do the problem? -6/2x(-1)?

OpenStudy (anonymous):

no...vertex is x=-b/2a...here b=-6 and a=-1. so: x=-(-6)/2(-1)=(6)/(-2)=-3

OpenStudy (anonymous):

ok, thank you...

OpenStudy (anonymous):

np

OpenStudy (anonymous):

here is your parabola's graph. your vertex is at x=-3 http://www.wolframalpha.com/input/?i=y%3D-x%5E2-6x

OpenStudy (anonymous):

hold the phone

OpenStudy (anonymous):

first coordinate of vertex is \(-3\) second coordinate is what you get when you replace \(x\) by \(-3\) you get \(-(-3)^2-6\times -3=-9+18=9\) so your second coordinate was correct, first one was wrong it is \((-3,9)\)

OpenStudy (anonymous):

yeah, I didn't bother with y :)